OpenStudy (anonymous):

Heres the problem: g(f(3)) g(x)=x−1 / 2 and f(x)=2x+1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Sorry guys. I solved this before and it didnt copy properly.

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

I'm solving it again.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

You would want to apply the f function and then the g fxn. So you start with f(x)=2x+1 and then make that x in g(x)=x−12 so g(f(x))=(2x+1)−12 and then you would substitute x for 3 and solve.

OpenStudy (anonymous):

The other way would be to solve through the process like so: 2x+1 --> 2(3)+1= 7 x=7 x-12 --> 7-12= g(f(x))= -5

OpenStudy (anonymous):

\[\frac{ 2(3) + 1 }{ 2 } - 1\] \[\frac{ 6 + 1 }{ 2 } - 1\] \[\frac{ 7 - 1 }{ 2 }\] \[\frac{ 6 }{ 2 }\] = 3

OpenStudy (anonymous):

Did I do something wrong?

OpenStudy (anonymous):

@Grazes ... I was wrong. Lol. Im going to compare answers now to see what I did wrong.

OpenStudy (anonymous):

It would help if you told me where you got \[\frac{ 2(3)+1 }{ 2 } -1\] from.

OpenStudy (anonymous):

@Grazes the entire question came out wrong because of an error on openstudy. Dont know why. Im going to retype the question here: Heres the actual QUESTION: Let f(x) = 2x + 1 and g(x) = x-1 / 2 find: g(f(3))

OpenStudy (anonymous):

Your answer is right.

OpenStudy (anonymous):

But your work is partially wrong. Check your first and 2nd expressions. the -1 should be in the fraction.

OpenStudy (anonymous):

Also, you should get into the habit of finding the g(f(x)) first because it helps a lot with harder problems with intensive calculation ((2x+1)-1)/2 2x/2=x x=3 and you're finished.

OpenStudy (anonymous):

oooooh !!! Hey @Grazes Im going to tag you in another. Can you tell me if I am correct when I do? Possibly I can put it below and tag you?

OpenStudy (anonymous):

If it's going to be in this post, you need not tag me, but sure.

OpenStudy (anonymous):

Problem: Let f(x) = 2x +1 \[g(x)=\frac{ x-1 }{ 2 }\] Find: f(g(x)) \[\frac{ 2(x-1) + 1}{ 2 }\] \[\frac{ 2x - 1 + 1 }{ 2 }\] \[\frac{ x - 1 + 1 }{ ? }\] \[x + 0\] \[= x\]

OpenStudy (anonymous):

Did I do that correctly?

OpenStudy (anonymous):

Yes. Also, I read the first problem incorrectly.

OpenStudy (anonymous):

The first one should have been \[2(\frac{ x-1 }{ 2 })+1\] \[(x-1)+1 = x\] The answer would be the same, though.

OpenStudy (anonymous):

Ohhhh, so I just need to work on organizing my problem easier for understanding and easy solving.

OpenStudy (anonymous):

Just don't get confused.

OpenStudy (anonymous):

Thanks so much ! @Grazes